OpenStudy (jfruscianterhcp):

find the number of diagonals of a regular polygon whose interior angle measures 144 degrees?

OpenStudy (jfruscianterhcp):

the answer is 35 but dunno how they get it

OpenStudy (jfruscianterhcp):

they only give the value of interior angle

OpenStudy (anonymous):

no of sides is 360/(exterior angle) =360/(180-44) =10 hence no of diagonals is 10*(10-3) /2 = 35

OpenStudy (dan815):



OpenStudy (anonymous):

irst find the number of sides and then for an sided poly gon the no of diagonals = n*(n-3)/2

OpenStudy (anonymous):

it is an n sided polygon. 180(n-2)=144n solve for n. what is n?

OpenStudy (primeralph):